OpenStudy (raffle_snaffle):

If a sum is worth $615 today and it will be worth $731 two years from now, how much was it worth one year ago?

OpenStudy (raffle_snaffle):

@tkhunny

OpenStudy (raffle_snaffle):

Okay so I have found the interest rate = 0.09023

OpenStudy (tkhunny):

Insufficient information. There are infinitely many pathways between the two values. Is there an interest accumulation specification?

OpenStudy (raffle_snaffle):

No this is all the info I have available.

OpenStudy (raffle_snaffle):

I have some charts in the back of my book that have compound interest factors available

OpenStudy (tkhunny):

You did this: 615(1+i)^2 = 731 ==> 9.0237537% Why not this: 615 + 615*r*2 = 731 ==> 9.4308943% Or this: 615(1+(I/12))^24 = 731 ==> 8.6707356% Or this: 615e^(2i) = 731 ==> 8.6395596% See what I mean? Worse yet, it could be a terribly unstable market - really good one year and horrible the next. 615 (1 + 0.25)(1 - 0.49105691) = 731 If all we're doing is annual compounding, then you have it. Seriously, INFINITELY many paths.

OpenStudy (raffle_snaffle):

I see.

OpenStudy (raffle_snaffle):

@mathstudent55